Prove that a.p of two numbers is greater than g.p..If and b are greater than 0.

Grade:10

1 Answers

sonika p
6 years ago
let a,b be two numbers, 
a.p = (a+b)/2
g.p = root ab
consider square of both a.p and g.p
a.p^2 = (a^2+b^2+2ab)/4 =  (a^2+b^2)/4+ab/2
g.p^2 = ab
a.p of two numbers is greater than g.p